OpenStudy (iwanttogotostanford):

Please help, I don't know how to do this and need to know how to: Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = Four divided by x squared. + 9

OpenStudy (iwanttogotostanford):

QUESTION: Find f(x) and g(x) so that the function can be described as y = f(g(x)). \[y=\frac{ 4 }{ x^2 } +9\]

OpenStudy (iwanttogotostanford):

@welshfella @Zyberg

OpenStudy (welshfella):

f(g(x) means 'function of g(x) ' where g(x) has replaced the x in the original f(x) so for example if f(x) = 3x^2 and g(x) = x - 1 then f(g(x)) = 3(x - 1)^2

OpenStudy (iwanttogotostanford):

so what would I do in my case?

OpenStudy (welshfella):

- so you have to sort f work backwards here

OpenStudy (iwanttogotostanford):

and how would i do that haha

OpenStudy (iwanttogotostanford):

@welshfella

OpenStudy (zyberg):

What are your guesses? ;) It would be easier to teach you, if we knew what you are thinking.

OpenStudy (welshfella):

suppose f(x) was x^2 + 9 what would g(x) have to be?

OpenStudy (welshfella):

sorry gtg right now

OpenStudy (iwanttogotostanford):

@Zyberg

OpenStudy (iwanttogotostanford):

really confused...

OpenStudy (zyberg):

I will say, that f(x) = x^2 + 9. Try to find g(x) ;)

OpenStudy (iwanttogotostanford):

what does g(x) even stand for again?

OpenStudy (zyberg):

A function :)

OpenStudy (iwanttogotostanford):

I know but what are the differences between f(x) and g(x). is it just that they are both different functions?

OpenStudy (iwanttogotostanford):

and how did you find that f(x) = x^ +9?? Where did the 4 go?

OpenStudy (zyberg):

You need to notice that 4 is 2^2 ;)

OpenStudy (zyberg):

Perhaps this would help you: http://www.purplemath.com/modules/fcncomp2.htm

OpenStudy (zyberg):

There are worked examples, so you might understand the concept better.

OpenStudy (iwanttogotostanford):

yes but then where does the "2" go.. if 4 is 2x^2

OpenStudy (iwanttogotostanford):

@phi can you help too?

OpenStudy (iwanttogotostanford):

yes But ii need to know how i can work THIS problem out, not a different one

OpenStudy (phi):

one way to do this is break your formula into two steps. can you do that ?

OpenStudy (zyberg):

Just read the examples in the provided link and try to understand them. That two and one x went to another function. You need to find 2 different functions from the one that is given.

OpenStudy (iwanttogotostanford):

I don't know how to start even on this problem except maybe plug in y for what f(x) is and then something for g(x)

OpenStudy (iwanttogotostanford):

thats why i came to open study, to get personalized help on this problem. the examples on the internet weren't helping much

OpenStudy (phi):

y= (4/x^2) + 9 you could break that into something + 9

OpenStudy (phi):

say g(x)= 4/x^2 and f(x)= x+9

OpenStudy (phi):

or g(x)= 2/x f(x) = x^2 + 9

OpenStudy (iwanttogotostanford):

so i could break it into 2^2

OpenStudy (phi):

g(x)= 4/x^2 -9 f(x) = x+18

OpenStudy (iwanttogotostanford):

ok! i get that part

OpenStudy (iwanttogotostanford):

these are my only choices though : f(x) = x + 9, g(x) = Four divided by x squared. f(x) = x, g(x) = Four divided by x. + 9 f(x) = One divided by x., g(x) = Four divided by x. + 9 f(x) = Four divided by x squared., g(x) = 9

OpenStudy (phi):

for each choice, in f(x)'s definition, replace x with g(x) then use g(x)'s definition to expand. one of the choices should match 4/x^2 + 9

OpenStudy (iwanttogotostanford):

i am confused what do you mean expand.? and what even is g(x)? Im sorry/// still a bit confused

OpenStudy (phi):

Here is an example. say g(x) = x+3 and f(x) = x/2 and we want to figure out f( g(x) ) in f(x)'s definition (i.e. x/2) look for any "x", and replace it with g(x) we get f( g(x) )= g(x) /2 (we replaced x with g(x) ) next , in that new definition, replace g(x) with it's definition i.e. with (x+3) in other words, in g(x)/2 we put in (x+3) instead of g(x) we get (x+3)/2 that is f(g(x))

OpenStudy (phi):

you could do that all in "one step", but it does not hurt to go slow until you get good at it.

OpenStudy (iwanttogotostanford):

oh ok, could you please do that but with this problem? I am creating a study guide and I already started o this problem and would like an example i already am working on

OpenStudy (phi):

try with the first choice.

OpenStudy (iwanttogotostanford):

so g(x)=2/x and then y= 4/x^2 +9 ?

OpenStudy (iwanttogotostanford):

I mean g(x)= 4/x^ and f(x)=x+9

OpenStudy (phi):

Here is the first choice f(x) = x + 9, g(x) = Four divided by x squared.

OpenStudy (phi):

yes, g(x)= 4/x^2 and f(x)= x+9 first step to finding f( g(x)) is to replace x in f's definition with g(x)

OpenStudy (iwanttogotostanford):

ok, so I would make it: f(x)= 4/x^2 + 9

OpenStudy (phi):

yes

OpenStudy (iwanttogotostanford):

ok, and then what would be my next step.. Thanks for you patience btw!

OpenStudy (phi):

your next step is to notice that is the answer

OpenStudy (iwanttogotostanford):

would it be this one: f(x) = Four divided by x squared., g(x) = 9 or this one: f(x) = x + 9, g(x) = Four divided by x squared. i think that its the last one

OpenStudy (phi):

you should follow the steps. f(x)= 4/x^2 , g(x)= 9 replace x in f's definition with g(x) we get f( g(x))= 4/( g(x) )^2 next, replace g(x) with its definition (which is just 9) f( g(x))= 4/9^2 = 4/81

OpenStudy (phi):

your first choice gives the formula you want: \[ f( g(x))= \frac{4}{x^2} + 9\]

OpenStudy (iwanttogotostanford):

thanks @phi